find the sum of the geometric series
1 + (1/2) + (1/4) + (1/8) + (1/16) + ....

Answer by stanbon(75887)

find the sum of the geometric series
1 + (1/2) + (1/4) + (1/8) + (1/16) + ....
-----------
It is a geometric series with r = (1/2)
Then
S(n) = 1/(1-r) = 1/(1/2) = 2
===============================
Cheers,
Stan H.